Analytical Geometry

Area of parallelogram formed by the lines y=mx, y=mx+1, y=nx and y=nx+1 equals

Hiteshu jani
7 Years agoGrade

1 Answer

Arun
7 Years ago
the two pairs of parallel lines pass through y=0 and y=1 
This distance on the y-axis forms the base of a triangle which is half the area of the parallelogram. 
The perpendicular height will be the distance between x=0 (y-axis) and the point of intersection of the lines y=mx and y=nx+1 (or the other pair)
y=mx . . . . . eqn 1 
y=nx+1. . . . eqn 2 
sub eqn 1 into eqn 2 
mx = nx +1 
mx - nx = 1 
x(m-n) = 1 
x = 1/ (m-n)
This is the x value and the distance the intersection is from the y-axis(perpendicular height)
Area = 2 * 1/2bh 
Area = bh 
Area = 1*1/(m-n) 
Area = 1/ (m-n)
as m-n could be positive or negative depending on the values of m and n, you need to take the absolute value so the area will be positive
giving 
Area = 1/ lm-nl